MAT 145: Topics In Contemporary Math
19: Solving Exponential Growth

3) A culture of paramecia contains 75 organisms. Assuming the presence of sufficient food
and physical room, the culture increases by 30% per hour. How long would it take for
there to be 288 paramecia? 288 because those unicellular organisms are too gross!

Answer by Boreal(15235)

288=75*1.4^t, t=hours
288/75=l.4^t
ln(288/75)=t ln (1.4)
t=4.00 hours
75*1.4^4=288.12, a para bit over two gross